# Mosaic: Composite Projection Pruning for Resource-efficient LLMs

### v1

@@ Abstract
Extensive compute and memory requirements limit the deployment of large language models (LLMs) on any hardware. Compression methods, such as pruning, can reduce model size, which in turn reduces resource requirements. State-of-the-art pruning is based on coarse-grained methods. They are time-consuming and inherently remove critical model parameters, adversely impacting the quality of the pruned model. This paper introduces projection pruning, a novel fine-grained method for pruning LLMs. In addition, LLM projection pruning is enhanced by a new approach we refer to as composite projection pruning - the synergistic combination of unstructured pruning that retains accuracy and structured pruning that reduces model size. We develop Mosaic, a novel system to create and deploy pruned LLMs using composite projection pruning. Mosaic is evaluated using a range of performance and quality metrics on multiple hardware platforms, LLMs, and datasets. Mosaic is 7.19 faster in producing models than existing approaches. Mosaic models achieve up to 84.2% lower perplexity and 31.4% higher accuracy than models obtained from coarse-grained pruning. Up to 67% faster inference and 68% lower GPU memory use is noted for Mosaic models.
Index Terms:
Composite projection pruning, Edge computing, Model compression, Large language models, Model pruning, Resource-Efficient LLMI Introduction
@@
Our work, Mosaic, is positioned to address these shortcomings. Existing pruning methods focus on coarse-grained pruning at the global and layer level of the LLM (further discussed in Section II). They prune every LLM component uniformly. This results in removing parts of the model that are critical to quality. Mosaic introduces novel fine-grained pruning of LLM projections. We leverage non-uniform pruning and apply it to different components of the LLM to selectively retain critical model parts. In addition, Mosaic synergistically combines unstructured and structured pruning in LLMs for the first time to create ‘composite projection pruning’. This pruning approach can produce compressed and resource-efficient LLMs that fit in limited memory and provide fast inference while having comparable quality to the foundation LLM. The models produced by Mosaic can be deployed on any hardware platform without requiring specific hardware/software accelerators.
@@
The remainder of this article is organized as follows. Section II presents the background for our work. Section III provides an overview of the methods underpinning Mosaic. Section IV presents the Mosaic modules. Section V evaluates Mosaic against relevant baselines. Section VI presents related work. Section VII concludes this article.
II Background and Motivation
@@ II-A LLM Architecture
II-B LLM Pruning Reduces Resource Footprint
@@ II-C Opportunities in LLM Pruning
In uniform pruning, every component (i.e. layer, block, and projection) of the LLM is pruned by the same amount, whereas in non-uniform pruning, some components are pruned more than others. Figure 3 shows the accuracy achieved and perplexity of uniform and non-uniform pruning on language benchmarks for the foundation LLM, LLaMa-3-8B. As more parameters are removed, non-uniform pruning achieves higher accuracy and lower perplexity than uniform pruning since non-uniform pruning can more accurately rank and prune LLMs. For the same accuracy loss, non-uniform pruning can eliminate up to 25% more parameters compared to uniform pruning, removing 70% versus 55% of the parameters, respectively. When 70% of parameters are removed, non-uniform pruning achieves a 40% lower perplexity of 25 compared to uniform pruning that has 40.
@@
Global Pruning is synonymous with uniform pruning. It was introduced to prune LLMs with lower GPU memory overheads [12]. Pruning occurs globally for a fixed percentage [21]. Each layer, block, and projection within the model removes the same percentage of parameters without accounting for their importance between different components.
Layer/Block Pruning is a more fine-grained pruning approach (we refer to as quasi-non-uniform pruning) in which each layer/block has a different pruning percentage [24]. For convolutional neural networks (CNNs), pruning layers by different percentages allows important parameters in sensitive layers to be untouched while aggressively pruning redundant layers [25]. This leads to better model accuracy at higher sparsities and extends to LLMs [22]. However, each projection within each layer/block is pruned to the same percentage, which does not account for the importance of parameters between the seven projections within each layer.
@@
II-D Our Work
We present Mosaic, a system that leverages the above opportunities and creates a novel approach for deriving and deploying compressed LLMs on resource-constrained hardware. Mosaic models bridge the performance and quality gap between UP and SP, allowing for flexible deployment of existing foundation LLMs to multiple target hardware platforms. While composite pruning is unique to Mosaic, the system can create pruned models using any of the three categories of model pruning depending on the available resources for the target platform. For example, a Mosaic LLM using UP will suit a GPU-rich environment to achieve high model quality. A Mosaic model derived using SP makes LLM inference possible at a minimal accuracy loss in environments that have limited resources. In addition, Mosaic models derived using composite projection pruning may be suited for weak GPUs and reducing the overall memory footprint.
III Mosaic Method
Mosaic adopts composite projection pruning on pre-trained foundation LLMs to create SLMs for a target device. Mosaic builds on the groundwork laid by global, layer, and block pruning while combining unstructured and structured pruning for the first time on LLMs and applies it for projections to achieve true non-uniform pruning. This section explores the proposed composite projection pruning method. When Mosaic is employed solely for unstructured or structured pruning, it utilizes the implementations outlined in Section V-A3.
III-A Pruning LLMs Across Projections
@@
III-B Composite Projection Pruning
@@
Figure 4 shows that for an LLM attention block, parameters in each projection are individually pruned via unstructured pruning based on POD (Section III-A) and simultaneously projection nodes between projections are pruned as a group via structured pruning. This combination is referred to as composite projection pruning and to derive SLMs from LLMs that are of high quality with low resource overheads. Section V-B3 will explore the benefits of composite projection pruning across a range of hardware.
@@ IV Mosaic System
Parameter Ranking Controller (RC) profiles the LLM to create a global rank () that represents the importance of each projection by identifying projection parameter outliers. Each LLM is profiled once to reuse the global rank for any pruning level across different pruned LLM variants. As a result, the initial overhead (Section V-B5) in creating the global rank is offset by faster inference achieved by the pruned LLMs.
@@ IV-A Implementation
Algorithm 1 is the Mosaic Parameter Ranking Controller (Figure 5). The RC generates the global rank using Equation 5 and Equation 6 to generate pruning targets for projections in the PC. First, the calibration samples , the to prune, and a constant for Equation 6 are loaded into system memory. The samples and LLM are then moved to the GPU memory (Line 1). Metadata, such as the number of layers and projections , are calculated by measuring the LLM dimensions (Lines 2-3). Lastly, an empty list of lists to store the global rank is initialized using and (Line 4). The CPU portion of the algorithm is then activated layer by layer as the samples are passed through the LLM (Lines 6-9). captures the activations (Line 8). For each projection, the weight metric is calculated by multiplying the norm of the activations by the magnitude of the projection weights (Equation 5, Line 11). The weight metric vector is flattened and concatenated across samples (Line 12). The mean of the weight metric is used with (Line 14) to calculate the ratio of projection outliers (Equation 6, Line 15) by dividing the number of outliers by the total parameters in the projection (Line 13). The global rank is then updated for the projection in the current layer (Line 16). The algorithm calculates the global rank for every projection across every layer . Finally, the normalized global rank is returned (Line 19).
V Experiments
@@
| P1 | AMD EPYC 7713 (x86) | 2 Nvidia A100 (80 GB, 1935 GB/s) |
| P2 | AMD EPYC 7713P (x86) | 2 Nvidia A6000 (48 GB, 768 GB/s) |
| P3 | Intel i9-13900KS (x86) | Nvidia RTX 3080 (10 GB, 760 GB/s) |
| P4 | Cortex-A78AE (ARM) | Nvidia AGX Orin (64 GB*, 205 GB/s) |
| P5 | Cortex-A76 (ARM) | Broadcom VideoCore VII (4 GB, 15 GB/s) |
@@
This section presents the experimental setup in Section V-A and the results from evaluation in Section V-B.
V-A Experimental Setup
V-A1 Testbed
Five hardware platforms as shown in Table I are used as the evaluation testbed: (P1) A 2 A100 GPU system akin to a Google Cloud a2-ultragpu-2g instance running Ubuntu 20.04.06 LTS with kernel version 5.4.0-193-generic is used for all experiments unless stated otherwise. (P2) A 2 A6000 GPU system akin to cloud GPU servers with lower GPU memory/bandwidth than P1. (P3) A consumer desktop GPU. (P4) A small form factor SoC GPU. (P5) A Raspberry Pi 5 with limited GPU memory and bandwidth.
V-A2 Evaluated Models
Five LLMs with varying characteristics shown in Table II are considered. Each model has seven projections per layer. All pruning methods are evaluated on each model for quality and performance metrics for a range of sparsity values. The LLMs have the following characteristics:
@@ V-A3 Pruning Methods
1) Global Pruning - each component is pruned uniformly by as presented in Section III-A.
@@
3) Composite Projection Pruning - Mosaic prunes parameters using unstructured pruning and then removes the lowest magnitude attention and feed-forward heads using structured pruning.
@@
V-A4 Dataset Metrics and Evaluation
11 natural language datasets across four tasks as shown in Table III are considered. They are summarized as follows:
@@ V-A4 Dataset Metrics and Evaluation
(3) Calibration - 128 samples from the C4 [40] dataset are used to calculate the activations in Equation 3 and Equation 5, and to calculate the inverse Hessian in SparseGPT.
@@ V-B Results
V-B1 Projection Pruning Performance (E1)
Global, layer and projection pruning is evaluated on all models in Table II on the perplexity and accuracy datasets in Table III. Each model is evaluated for varying sparsities up to 80% as the model collapses beyond this range [12, 42]. Perplexity is reported per dataset, whereas accuracy is the mean across the seven datasets shown in Table III.
@@ V-B1 Projection Pruning Performance (E1)
Accuracy: Table IV shows the mean zero-shot accuracy of LLaMa-3.1-8B and LLaMa-2-13B for each pruning method at 20%, 40%, 60%, and 80% sparsity. 0% sparsity is the accuracy of the original LLM with no pruning. Projection pruning maintains the highest accuracy for both models and all sparsity values. At 40% sparsity, projection pruning improves accuracy by less than 1% for LLaMa-2-13B against other methods. However, at 80% sparsity, this difference increases to 31.4% higher accuracy. Similarly, LLaMa-3.1-8B achieves 13.2% higher accuracy at the same sparsity level.
@@
V-B2 Projection Pruning Control (E2)
This experiment examines the empirical differences between the granularity of global, layer and projection pruning. Figure 8 shows the pruning target of each layer of LLaMa-3.1-8B pruned by 80% for each method. On WikiText-2, global, layer and projection pruning achieve a perplexity of 221, 182 and 82, respectively. On PTB, 320, 306 and 166, respectively.
@@ V-B2 Projection Pruning Control (E2)
V-B3 Composite Projection Pruning Performance (E3)
Composite projection pruning (Figure 4) is evaluated against unstructured and structured projection pruning. Figure 9 shows inference latency and memory usage of LLaMa-7B on the five hardware platforms shown in Table I for varying pruning targets. Inference is carried out with 2048 tokens input and an output of 128 tokens per MLPerf standardized benchmark [43] with a batch size of 12. On P5, input tokens of 128 and output tokens of 16 with a batch size of 1 are used due to memory and compute constraints. Inference latency is the wall clock time for processing the entire input string and generating the output string. GPU memory used on each platform includes model weights, activations, attention, the software libraries and the Mosaic framework. The size of software libraries, such as Python packages compiled for different architectures, varies across platforms, affecting the size of memory allocated. On P3 and P5 with low GPU memory, the model weights are offloaded to device storage using swap memory when exceeding GPU memory capacity (dashed gray lines in Figure 9). Table V shows the perplexity achieved on LLaMa-7B for different pruning methods.
System Performance: Composite and structured pruning reduce inference time and GPU memory used as more parameters are removed. Unstructured pruning does not offer any performance benefits. Although structured pruning has better runtime performance, model quality measured by perplexity (Table V) is impacted. When 80% of the parameters are removed, composite pruning has 30% - 67% lower inference latency than unstructured pruning while reducing memory use by 60% - 68%.
Low Memory Platforms: The GPU memory available on P3 and P5 is lower than the foundation model of LLaMa-7B. Therefore, model layers are offloaded to storage, and only a subset is loaded into memory. During inference, layers are transferred in and out of memory and storage, resulting in a data transfer bottleneck, thus increasing inference latency. For example, on P3, inference latency is 420s when using offloading but once a model is pruned enough to use less than 10GB of GPU memory, inference latency is reduced by up to 30 (14s for 80% pruned Mosaic model). The foundation model and the pruned models obtained from unstructured pruning cannot be run on P5. When the memory required is reduced from 10 GB to 7.9 GB using Mosaic, the inference latency reduces from 400s to 110s.

### v2

@@ Abstract
Extensive compute and memory requirements limit the deployment of large language models (LLMs) on any hardware. Compression methods, such as pruning, can reduce model size, which in turn reduces resource requirements. State-of-the-art pruning is based on coarse-grained methods. They are time-consuming and inherently remove critical model parameters, adversely impacting the quality of the pruned model. This paper introduces projection pruning, a novel fine-grained method for pruning LLMs. In addition, LLM projection pruning is enhanced by a new approach we refer to as composite projection pruning - the synergistic combination of unstructured pruning that retains accuracy and structured pruning that reduces model size. We develop Mosaic, a novel system to create and deploy pruned LLMs using composite projection pruning. Mosaic is evaluated using a range of performance and quality metrics on multiple hardware platforms, LLMs, and datasets. Mosaic is 7.19 faster in producing models than existing approaches. Mosaic models achieve up to 84.2% lower perplexity and 31.4% higher accuracy than models obtained from coarse-grained pruning. Up to 67% faster inference and 68% lower GPU memory use is noted for Mosaic models. Mosaic is available for public use from https://github.com/blessonvar/Mosaic
keywords:
Composite projection pruning , Edge computing , Model compression , Large language models , Model pruning , Resource-efficient LLM[sta]organization=School of Computer Science, University of St Andrews,addressline=Jack Cole Building, city=St Andrews, postcode=KY16 9SX, state=Fife, country=Scotland, United Kingdom
[rmi]organization=Autonomous Networking Research & Innovation Department, Rakuten Mobile, Inc.,addressline=Rakuten Crimson House, 1-14-1 Tamagawa, Setagaya-ku, city=Tokyo, postcode=158-0094, state=Tokyo, country=Japan
1 Introduction
@@
Our work, Mosaic, is positioned to address these shortcomings. Existing pruning methods focus on coarse-grained pruning at the global and layer level of the LLM (further discussed in Section 2). They prune every LLM component uniformly. This results in removing parts of the model that are critical to quality. Mosaic introduces novel fine-grained pruning of LLM projections. We leverage non-uniform pruning and apply it to different components of the LLM to selectively retain critical model parts. In addition, Mosaic synergistically combines unstructured and structured pruning in LLMs for the first time to create ‘composite projection pruning’. This pruning approach can produce compressed and resource-efficient LLMs that fit in limited memory and provide fast inference while having comparable quality to the foundation LLM. The models produced by Mosaic can be deployed on any hardware platform without requiring specific hardware/software accelerators.
@@
The remainder of this article is organized as follows. Section 2 presents the background for our work. Section 3 provides an overview of the methods underpinning Mosaic and the design of the Mosaic modules. Section 4 evaluates Mosaic against relevant baselines. Section 5 presents related work. Section 6 concludes this article.
2 Background and Motivation
@@ 2.1 LLM Architecture
2.2 LLM Pruning Reduces Resource Footprint
@@ 2.3 Opportunities in LLM Pruning
In uniform pruning, every component (i.e. layer, block, and projection) of the LLM is pruned by the same amount, whereas in non-uniform pruning, some components are pruned more than others. Figure 3 shows the accuracy and perplexity achieved by uniform and non-uniform pruning on language benchmarks for the foundation LLM, LLaMa-3-8B. As more parameters are removed, non-uniform pruning achieves higher accuracy and lower perplexity than uniform pruning since non-uniform pruning can more accurately rank and prune LLMs. For the same accuracy loss, non-uniform pruning can eliminate up to 25% more parameters compared to uniform pruning, removing 70% versus 55% of the parameters, respectively. When 70% of parameters are removed, non-uniform pruning achieves a 40% lower perplexity of 25 compared to uniform pruning that has 40.
@@
Global Pruning is synonymous with uniform pruning. It was introduced to prune LLMs with lower GPU memory overheads [12]. Pruning occurs globally for a fixed percentage [21]. Each layer, block, and projection within the model removes the same percentage of parameters without accounting for their importance between different components. While projections are pruned, they are not evaluated individually for importance; instead, they are uniformly pruned by applying the same percentage across all projections.
Layer/Block-wise Pruning is a more fine-grained pruning approach (we refer to as quasi-non-uniform layer/block pruning) in which each layer/block has a different pruning percentage [24]. For convolutional neural networks (CNNs), pruning layers by different percentages allows important parameters in sensitive layers to be untouched while aggressively pruning redundant layers [25]. This leads to better model accuracy at higher sparsities and extends to LLMs [22]. However, each projection within each layer/block is pruned to the same percentage, which does not account for the importance of parameters between the seven projections within each layer.
@@
Global/Layer/Projection pruning alone refers to the granularity of the pruning method, which can be applied in an unstructured or structured manner. Therefore, the potential dimensions of projection pruning that are explored in this manuscript include unstructured, structured, and composite projection pruning.
2.4 Leveraging the Opportunities
We present Mosaic, a system that leverages the above opportunities and creates a novel approach for deriving and deploying compressed LLMs on resource-constrained hardware. Mosaic models bridge the performance and quality gap between UP and SP, allowing for flexible deployment of existing foundation LLMs to multiple target hardware platforms. While composite projection pruning is unique to Mosaic, the system can create pruned models using any of the three categories of model pruning depending on the available resources for the target platform. For example, a Mosaic LLM using UP will suit a GPU-rich environment to achieve high model quality. A Mosaic model derived using SP makes LLM inference possible at a minimal accuracy loss in environments that have limited resources. In addition, Mosaic models derived using composite projection pruning may be suited for weak GPUs and reducing the overall memory footprint.
3 Design of Mosaic
Mosaic adopts composite projection pruning on pre-trained foundation LLMs to create SLMs for a target device. Mosaic builds on the groundwork laid by global, layer, and block pruning while combining unstructured and structured pruning for the first time on LLMs and applies it for projections to achieve true non-uniform pruning. This section explores the proposed composite projection pruning method. When Mosaic is employed solely for unstructured or structured pruning, it utilizes the implementations outlined in Section 4.1.3.
3.1 Pruning LLMs Across Projections
@@
3.2 Composite Projection Pruning
@@
Figure 4 shows that for an LLM attention block, parameters in each projection are individually pruned via unstructured pruning based on POD (Section 3.1) and simultaneously projection nodes between projections are pruned as a group via structured pruning. This combination is referred to as composite projection pruning and to derive SLMs from LLMs that are of high quality with low resource overheads. Section 4.2.3 will explore the benefits of composite projection pruning across a range of hardware.
@@ 3.3 Mosaic System
Parameter Ranking Controller (RC) profiles the LLM to create a global rank () that represents the importance of each projection by identifying projection parameter outliers. Each LLM is profiled once to reuse the global rank for any pruning level across different pruned LLM variants. As a result, the initial overhead (Section 4.2.5) in creating the global rank is offset by faster inference achieved by the pruned LLMs.
@@ 3.4 Implementation
Algorithm 1 shows the procedure of the Mosaic Parameter Ranking Controller (Figure 5). The RC generates the global rank using Equation 5 and Equation 6 to generate pruning targets for projections in the PC. First, the calibration samples , the to prune, and a constant for Equation 6 are loaded into system memory. The samples and LLM are then moved to the GPU memory (Line 1). Metadata, such as the number of layers and projections , are calculated by measuring the LLM dimensions (Lines 2-3). Lastly, an empty list of lists to store the global rank is initialized using and (Line 4). The CPU portion of the algorithm is then activated layer by layer as the samples are passed through the LLM (Lines 6-9). captures the activations (Line 8). For each projection, the weight metric is calculated by multiplying the norm of the activations by the magnitude of the projection weights (Equation 5, Line 11). The weight metric vector is flattened and concatenated across samples (Line 12). The mean of the weight metric is used with (Line 14) to calculate the ratio of projection outliers (Equation 6, Line 15) by dividing the number of outliers by the total parameters in the projection (Line 13). The global rank is then updated for the projection in the current layer (Line 16). The algorithm calculates the global rank for every projection across every layer . Finally, the normalized global rank is returned (Line 19).
4 Experiments
@@
[…]
@@
This section presents the experimental setup in Section 4.1 and the results from evaluation in Section 4.2.
4.1 Experimental Setup
4.1.1 Testbed
Five hardware platforms as shown in Table 1 are used as the evaluation testbed: (P1) A 2 A100 GPU system akin to a Google Cloud a2-ultragpu-2g instance running Ubuntu 20.04.06 LTS with kernel version 5.4.0-193-generic is used for all experiments unless stated otherwise. (P2) A 2 A6000 GPU system akin to cloud GPU servers with lower GPU memory/bandwidth than P1. (P3) A consumer desktop GPU. (P4) A small form factor SoC GPU. (P5) A Raspberry Pi 5 with limited GPU memory and bandwidth.
4.1.2 Evaluated Models
Five LLMs with varying characteristics shown in Table 2 are considered. Each model has seven projections per layer. All pruning methods are evaluated on each model for quality and performance metrics for a range of sparsity values. The LLMs have the following characteristics:
@@ 4.1.3 Pruning Methods
1) Global Pruning - each component is pruned uniformly by as presented in Section 3.1.
@@
3) Composite Projection Pruning - Mosaic first applies unstructured pruning to individual parameters, followed by structured pruning to remove the lowest-magnitude attention and feed-forward heads. This approach is analogous to prior composite pruning methods developed for CNNs [26].
@@
4.1.4 Dataset Metrics and Evaluation
11 natural language datasets across four tasks as shown in Table 3 are considered. These datasets cover a wide range of dynamic inputs and are used to evaluate model performance across diverse domains and task settings. They are summarized as follows:
@@ 4.1.4 Dataset Metrics and Evaluation
(3) Calibration - As considered in the existing pruning literature, 128 samples from the C4 [40] dataset are used to calculate the activations in Equation 3 and Equation 5, and to calculate the inverse Hessian in SparseGPT.
@@ 4.2 Results
4.2.1 Projection Pruning Performance (E1)
Global, layer and projection pruning is evaluated on all models in Table 2 on the perplexity and accuracy datasets in Table 3. Each model is evaluated for varying sparsities up to 80% as the model collapses beyond this range [12, 42]. Perplexity is reported per dataset, whereas accuracy is the mean across the seven datasets shown in Table 3.
@@ 4.2.1 Projection Pruning Performance (E1)
Accuracy: Table 4 shows the mean zero-shot accuracy of LLaMa-3.1-8B and LLaMa-2-13B for each pruning method at 20%, 40%, 60%, and 80% sparsity. 0% sparsity is the accuracy of the original LLM with no pruning. Projection pruning maintains the highest accuracy for both models and all sparsity values. At 40% sparsity, projection pruning improves accuracy by less than 1% for LLaMa-2-13B against other methods. However, at 80% sparsity, this difference increases to 31.4% higher accuracy. Similarly, LLaMa-3.1-8B achieves 13.2% higher accuracy at the same sparsity level.
@@
4.2.2 Projection Pruning Control (E2)
This experiment examines the empirical differences between global, layer, and projection pruning, highlighting how the choice of granularity significantly affects pruning outcomes. Figure 8 shows the pruning target of each layer of LLaMa-3.1-8B pruned by 80% for each method. On WikiText-2, global, layer and projection pruning achieve a perplexity of 221, 182 and 82, respectively. On PTB, 320, 306 and 166, respectively.
@@ 4.2.2 Projection Pruning Control (E2)
4.2.3 Composite Projection Pruning Performance (E3)
This experiment compares composite projection pruning against unstructured and structured baselines across a range of performance and accuracy metrics. Figure 9 shows inference latency and memory usage of LLaMa-7B on the five hardware platforms shown in Table 1 for varying pruning targets. Inference is carried out with 2048 tokens input and an output of 128 tokens per MLPerf standardized benchmark [43] with a batch size of 12. On P5, input tokens of 128 and output tokens of 16 with a batch size of 1 are used due to memory and compute constraints. Inference latency is the wall clock time for processing the entire input string and generating the output string. GPU memory used on each platform includes model weights, activations, attention, the software libraries and the Mosaic framework. The size of software libraries, such as Python packages compiled for different architectures, varies across platforms, affecting the size of memory allocated. On P3 and P5 with low GPU memory, the model weights are offloaded to device storage using swap memory when exceeding GPU memory capacity (dashed gray lines in Figure 9). Table 5 shows the perplexity achieved on LLaMa-7B for different pruning methods.
System Performance: Composite and structured pruning reduce inference time and GPU memory used as more parameters are removed. Unstructured pruning does not offer any performance benefits. Although structured pruning has better runtime performance, model quality measured by perplexity (Table 5) is impacted. When 80% of the parameters are removed, composite pruning has 30% - 67% lower inference latency than unstructured pruning while reducing memory use by 60% - 68%.
Low Memory Platforms: The GPU memory available on P3 and P5 is lower than the foundation model of LLaMa-7B. Therefore, model layers are offloaded to storage, and only a subset is loaded into memory. During inference, layers are transferred in and out of memory and storage, resulting in a data transfer bottleneck, thus increasing inference latency. For example, on P3, inference latency is 420s when using offloading but once a model is pruned enough to use less than 10GB of GPU memory, inference latency is reduced by up to 30 (14s for 80% pruned Mosaic model). The foundation model and the pruned models obtained from unstructured pruning cannot be run on P5. When the memory required is reduced from 10 GB to 7.9 GB using Mosaic, the inference latency reduces from 400s to 110s. Composite pruning of Mosaic enables model deployment on memory-constrained hardware by structurally reducing the memory footprint, overcoming the limitations of unstructured pruning, which alone does not reduce inference latency or enable execution under tight memory budgets. This approach allows Mosaic to adapt the sparsification strategy to hardware constraints by applying structured pruning where unstructured sparsity alone cannot be exploited due to a lack of hardware accelerators or limited memory capacity.